OpenStudy (anonymous):

Alright I was helping my friend with some high school math and she had the following question: if 2 + sqrt(3) is a polynomial root, name another root of the polynomial, and explain how you know it must also be a root

OpenStudy (anonymous):

I know the answer should be: \[2 - \sqrt3\] but here is the thing.... the question never gives a polynomial. It just gives a condition that the polynomial must statisfy. Well doesn't \[(x -2 - \sqrt3)^2\] satisfy that ciondition? Yet the only root of this polynomial is \[2 + \sqrt3\]

OpenStudy (anonymous):

so basically I managed to confuse myself on something basic...I know its something I am forgetting but I don't know what.

OpenStudy (anonymous):

the original question isn't 2+i*sqrt(3) and all coef are real?

OpenStudy (anonymous):

http://imgur.com/A4D2BYl

hero (hero):

The basis for the question itself comes from a property of radical conjugate roots. If a polynomial \(P(x)\) with rational coefficients has \(a + \sqrt{b}\) as a root, where a, b are rational and \(\sqrt{b}\) is irrational, then \(a − \sqrt{b}\) is also a root. http://en.wikipedia.org/wiki/Properties_of_polynomial_roots#Radical_conjugate_roots

OpenStudy (anonymous):

I found that also which is why I know what the answer should be but I still can't see what is wrong with my possible polynomial

OpenStudy (anonymous):

There has to be something wrong with my proposition because it seems to defy the property of radical conjugate roots.

OpenStudy (anonymous):

@Hero but he dont mention taht the coef are rational

OpenStudy (anonymous):

lol at least I'm not the only one this is a problem for lol

hero (hero):

The product of \((x - 2 - \sqrt{3})^2\) produces co-efficients that are not rational.

OpenStudy (anonymous):

...so a polynomial has to have rational coefficients?

hero (hero):

Technically, since the question itself doesn't mention anything about rational roots, your response could also be accepted, but you should present your argument in the following manner: Assume \(x = 2 + \sqrt{3}\) Then \(x - 2 - \sqrt{3} = 0\) and \((x - 2 - \sqrt{3})^2 = 0\) yet the graph of \(y = (x - 2 - \sqrt{3})^2 \) produces only one root.

hero (hero):

*rational coefficients not rational roots

OpenStudy (anonymous):

....but in the wiki you linked it says nothing about the coefficients being rational

OpenStudy (anonymous):

it says the coefficents are members of the field of complex numbers which, excuse me if I'm wrong, the coefficents of my polynomial are

OpenStudy (anonymous):

@Hero please answer, I'm still a little confused

hero (hero):

I'll link you to it once more. You have to align your eyes to the correct area on the page. Look for the words "rational coefficients" while scanning the page. http://en.wikipedia.org/wiki/Properties_of_polynomial_roots#Radical_conjugate_roots

OpenStudy (anonymous):

...oops lol I'll blame the fact that the page said this on the first line: where the ai belong to some field, which, in this article, is always the field \mathbb C of the complex numbers

OpenStudy (anonymous):

I get it now thanks :P

hero (hero):

I think you were looking at the wrong area. I linked you a specific area on the page.

hero (hero):

The part that is entitled "Radical Conjugate Roots"

OpenStudy (anonymous):

Yup I was. Thanks again. Just a poorly written question on her instructor's part (not surprising with this guy). I just didn't see that this time.